OpenStudy (anonymous):

60% of what number is 72? 35% of $200 is how much? 12 is what percent of 20? 12% of 20 is? 10% of what number is 71? what percent of 50 is 15? 16 is what percent of 80? please help

OpenStudy (anonymous):

Hint X x 60/100 = 72. Find X.

OpenStudy (jackellyn):

20*0.12=? 200*0.35=?